OpenStudy (anonymous):

the sum of two numbers is 60. five times the smaller number minus twice the larger number is 223. what are the number

OpenStudy (anonymous):

x+y=60 5x-2y=223 Then solve for x and y

OpenStudy (anonymous):

let the number be x and 60-x as per the question we have 5x-2(60-x)=223 5x -120 +2x =223 7x =223+120 or 7x= 343 or x=343/7 hence x=49 thus the numbers 49 and 11

OpenStudy (anonymous):

is the quetion : the sum of two numbers is 60. five times the larger number minus twice the smaller number is 223. what are the number

OpenStudy (anonymous):

You got the right answer @matrick. Smaller number is 11 and bigger one is 49

OpenStudy (anonymous):

thanks but the question must be synonymus....

OpenStudy (anonymous):

Let x be the smaller number Let y be the larger number x < y x + y = 60 5x - 2y = 223 60 - y < y 60 < 2y 30 < y x < 30 Now what...? 2x + 2y = 120 5x - 2y = 223 7x = 343 x = 49 but this can't be true, hmm...? This question doesn't make sense